3. Two identical size, flat, disc, conductors, are oriented parallel to each other. Each plate has a radius of 0.150m. respecPlease do problem 3 and problem 4 .
Answer #1

3 Area of each plate 2A = 22 A = 71 (0.15)² = 0.0708m 2 V = 50v Ve=15V - Uc d plate ① plate & E= AV = V, -va d V, - V2 = d 50Ausy) q = 50x10-12 c. 22 - 22.58 10-12 c. Ri= 10x10 3 m; R2 = 15:08 10 0 Kq , + K92 V, = K ( a + r) V = 9x 109/50x10 29.5-x1